OpenStudy (anonymous):

I've got 4 algebra 2 questions that I'm really stuck on. can someone please help me??

OpenStudy (anonymous):

i've answered number 2, but I'm not sure about that answer.

OpenStudy (anonymous):

@Michele_Laino @nincompoop @campbell_st @Loser66

OpenStudy (anonymous):

@Michele_Laino will you help me?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

exercise #1

OpenStudy (michele_laino):

here we have to apply the Theore of carnot, and we can write: \[\Large Q{R^2} = P{Q^2} + P{R^2} - 2PQ \times PR\cos P\]

OpenStudy (michele_laino):

substituting your data, we get: \[\Large QR = \sqrt {{{7.5}^2} + {{8.4}^2} - 27.5 \times 8.4 \times \cos 43} = ...?\] what is QR?

OpenStudy (michele_laino):

oops... \[\Large QR = \sqrt {{{7.5}^2} + {{8.4}^2} - 2 \times 7.5 \times 8.4 \times \cos 43} = ...?\] what is QR?

OpenStudy (anonymous):

sqrt −231cs64+126.81

OpenStudy (anonymous):

find QR and then use sin formula to find <R

OpenStudy (anonymous):

I don't know how to do that..

OpenStudy (michele_laino):

I got this: QR= 5.89

OpenStudy (anonymous):

oh..how did I get that answer....??

OpenStudy (michele_laino):

next we have to apply the law of sines, so we can write: \[\Large \frac{{QR}}{{\sin P}} = \frac{{PQ}}{{\sin R}}\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

substituting our data, we get: \[\Large \sin R = \frac{{PQ}}{{QR}}\sin P = \frac{{7.5}}{{5.89}} \times \sin 43 = ...?\]

OpenStudy (anonymous):

0.86841896

OpenStudy (michele_laino):

ok! so what is R?

OpenStudy (anonymous):

0.86841896 <----this?

OpenStudy (michele_laino):

no, that is sin(R)

OpenStudy (anonymous):

I'm not sure then..

OpenStudy (michele_laino):

you have to use the sin^-1 function on your calculator

OpenStudy (michele_laino):

I got this: R= 60.27 degrees

OpenStudy (anonymous):

Oh, now I see. So it is D?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thank you!!! can you help me on the others?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

for the second one, I'm not sure if that one is the correct answer.

OpenStudy (michele_laino):

exercise #2

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

we have: \[\Large C = 180 - \left( {A + B} \right) = 180 - \left( {41 + 32} \right) = ...?\]

OpenStudy (michele_laino):

what is C?

OpenStudy (anonymous):

107

OpenStudy (anonymous):

\[\frac{\text{AB}}{\sin ((180-32-41) {}^{\circ})}=\frac{9}{\sin (41 {}^{\circ})} \]\[\text{AB}\to 13.1189 \]

OpenStudy (anonymous):

Is it A then?

OpenStudy (michele_laino):

ok! now we have to apply the law of sines again, so we can write: \[\Large \frac{{AB}}{{\sin C}} = \frac{{AC}}{{\sin B}}\]

OpenStudy (anonymous):

@robtobey did this part of the problem already, right?

OpenStudy (anonymous):

A.

OpenStudy (michele_laino):

substituting your data, we get: \[\Large AB = \frac{{AC\sin C}}{{\sin B}} = \frac{{9 \times \sin 107}}{{\sin 32}} = ...?\]

OpenStudy (anonymous):

13.1189

OpenStudy (michele_laino):

I got 16.24

OpenStudy (anonymous):

I was using @robtobey 's answer. I just did the calculations, I got 16.24 also

OpenStudy (anonymous):

It's C

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

I get that one too...thank you!!

OpenStudy (anonymous):

This is really helpful

OpenStudy (michele_laino):

now let's go to exercise #3

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

here we have to apply the subsequent formula: \[\Large \begin{gathered} Area = \frac{1}{2} \times PQ \times RQ \times \sin Q = \hfill \\ \hfill \\ = \frac{1}{2} \times 18 \times 8 \times \sin 25 =...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

30.4

OpenStudy (michele_laino):

yes! I got Area=30.42

OpenStudy (anonymous):

yay!!

OpenStudy (michele_laino):

Next, let's go on exercise #4

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

if: \[\Large A + B = C\] then: \[\Large B = C - A\] am I right?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

so we can write this: \[\Large \begin{gathered} B = C - A = \left( {\begin{array}{*{20}{c}} 1&{ - 2}&8 \\ 5&0&2 \end{array}} \right) - \left( {\begin{array}{*{20}{c}} 4&2&7 \\ 3&7&{ - 3} \end{array}} \right) = \hfill \\ \hfill \\ = \left( {\begin{array}{*{20}{c}} {1 - 4}&{...}&{...} \\ {...}&{...}&{...} \end{array}} \right) \hfill \\ \end{gathered} \] please complete

OpenStudy (anonymous):

1 -4 -3 2 -7 5

OpenStudy (anonymous):

Very sorry for my mistake. Labeled BC and AC as both 9 each. 16.24 is correct. I believe that this is the first or second of a problem error or mine discovered by the other participants.

OpenStudy (anonymous):

THat's okay, we all make mistakes :)

OpenStudy (michele_laino):

I meant 1-4= -3

OpenStudy (anonymous):

What?

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} B = C - A = \left( {\begin{array}{*{20}{c}} 1&{ - 2}&8 \\ 5&0&2 \end{array}} \right) - \left( {\begin{array}{*{20}{c}} 4&2&7 \\ 3&7&{ - 3} \end{array}} \right) = \hfill \\ \hfill \\ = \left( {\begin{array}{*{20}{c}} {1 - 4}&{...}&{...} \\ {...}&{...}&{...} \end{array}} \right) = \left( {\begin{array}{*{20}{c}} { - 3}&{...}&{...} \\ {...}&{...}&{...} \end{array}} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ooh so what do I do next?

OpenStudy (michele_laino):

you have to subtract every component of the second matrix, from the corresponding component of the first matrix

OpenStudy (anonymous):

-3 0 1 2 7 1

OpenStudy (michele_laino):

for example: \[\Large \begin{gathered} B = C - A = \left( {\begin{array}{*{20}{c}} 1&{ - 2}&8 \\ 5&0&2 \end{array}} \right) - \left( {\begin{array}{*{20}{c}} 4&2&7 \\ 3&7&{ - 3} \end{array}} \right) = \hfill \\ \hfill \\ = \left( {\begin{array}{*{20}{c}} {\left( {1 - 4} \right)}&{\left( { - 2 - 2} \right)}&{\left( {8 - 7} \right)} \\ {\left( {5 - 3} \right)}&{\left( {0 - 7} \right)}&{\left( {2 - \left( { - 3} \right)} \right)} \end{array}} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

mine wasn't right?

OpenStudy (michele_laino):

no, I'm sorry

OpenStudy (anonymous):

Wait, I'm onfused. How do I do this?

OpenStudy (michele_laino):

it is simple: 1-4= -3 -2-2=-4 and so on...

OpenStudy (anonymous):

-3 -4 -1 2 7 -1

OpenStudy (anonymous):

oh, no the 1's aren't negatives.

OpenStudy (michele_laino):

0-7= -7 right?

OpenStudy (michele_laino):

2-(-3)= 2+ 3= 5

OpenStudy (michele_laino):

8-7=1 and, finally: 5-3=2 \[\Large \begin{gathered} B = C - A = \left( {\begin{array}{*{20}{c}} 1&{ - 2}&8 \\ 5&0&2 \end{array}} \right) - \left( {\begin{array}{*{20}{c}} 4&2&7 \\ 3&7&{ - 3} \end{array}} \right) = \hfill \\ \hfill \\ = \left( {\begin{array}{*{20}{c}} {\left( {1 - 4} \right)}&{\left( { - 2 - 2} \right)}&{\left( {8 - 7} \right)} \\ {\left( {5 - 3} \right)}&{\left( {0 - 7} \right)}&{\left( {2 - \left( { - 3} \right)} \right)} \end{array}} \right) = \hfill \\ \hfill \\ = \left( {\begin{array}{*{20}{c}} { - 3}&{ - 4}&1 \\ 2&{ - 7}&5 \end{array}} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Oh, I see what I was doing wrong now.

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

A

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thank you so much for your help!!

OpenStudy (michele_laino):

:)